[Brahms-l] Shift report 20040109 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Fri Jan 09 2004 - 00:15:00 EST
------------------------------------------------------------
Date: Fri Jan  9 00:14:59 2004, Shift: 16:00-24:00
Supervisor: JHL, with: 
Log book#15 pages 26 - 26

Shift Summary:
------------------------------------------------------------
 Beam stored for ~3hrs:
 -ZDC 1200 -> 500/sec 
 -ZDC/TOFW: 1/1.4 with and without collimator

 Data:
 -FS 4deg 1B ~900K trigger6 (runs 9483,9484)
  (MRS at 90deg without TPM2)

 Gas flow seems all a lit low: adjusted
 Trigger2 being setup by fv
 No new detector problems
 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 9 00:15:16 2004

This archive was generated by hypermail 2.1.8 : Fri Jan 09 2004 - 00:15:32 EST